How many ways can you order a hamburger if you can order it with or without the following items: cheese, lettuce, tomatoes, pickles or onions?
Found 3 solutions by stanbon, JWG, richard1234:
Answer by stanbon(48568)

How many ways can you order a hamburger if you can order it with or without the following items: cheese, lettuce, tomatoes, pickles or onions?
You choose to have or you choose to not have. 2 choices
But you do this 5 times
# of orders = 2^5=32 different orders.
Cheers,
Stan H.

Answer by JWG(21)

My solution is much longer, but same answer.
I read this problem as saying, "How many ways can you make a hamburger with five different fixings?" This means I can make a hamburger with all five fixings, or as little as none--just a plain burger with a bun.
Five Fixings: 1 way
Four Fixings: 5C4 = 5 ways
Three Fixings: 5C3 = 10 ways
Two Fixings: 5C2 = 10 ways
One Fixing: 5 ways
Zero Fixings: 1 way

1+5+10+10+5+1 = 32
32 ways to make a hamburger when there are five fixings to choose from, and amount of the five chosen vary.

Answer by richard1234(4802)

Each item can be put on or left on. There are five items, so there are 2^5, or 32, possible combinations of the burger.

Another way to visualize it is to note that all of the binary numbers from 00000 to 11111 represent the numbers 0 through 31, in which there are 32 numbers here.